Andrew Marzo

Content and Structure – The paper

1. (10) Has a clearly stated topic or question that it will address – 10

2. (5) Does address the stated topic or question – 5

3. Has a clearly articulated and logical structure:

a. (10) First paragraph lays out the topic or question – 10 b. (10) Later paragraphs have topic sentences and clearly identifiable main ideas –

9 c. (10) Transitions between paragraphs are logical – 9

4. (10) Accurately presents facts and viewpoints of sources – 8

5. (10) Provides sufficient textual evidence to support the argument – 10

6. (10) Is concise, clear, and not too repetitive – 9

7. (10) Has a clearly identifiable conclusion – 10

8. (15) Has a convincing argument that draws together the evidence presented – 10

Total: 90 (A-)

Style and English Usage – The paper

1. (20) Is clear, direct and appropriate in tone – 18

2. (10) Is free of misspellings – 10

3. (15) Is free of grammatical mistakes (e.g. subject/verb agreement, proper use of pronouns) – 13

4. (10) Lacks incomplete sentences – 10

5. (10) Lacks run-on or overly complicated sentences – 10

6. (10) Uses correct punctuation – 7

7. (15) Is free of jargon, cliches, and misused words – 13

8. (10) Cites quotations and/or references correctly – 10

Total: 91 (A-)

Comments:

1. I think the issue was not just their expertise (their knowledge of their field), but also the attitudes of the patients. I think patients used to have a sort of awe of doctors that made it unthinkable that doctors would be questioned in many cases.

2. There is some truth to this. But then why do you suppose that Skloot included the stories of the much more drastic mistreatment of African Americans in the Tuskegee study and similar treatment of other people who could not defend their own rights in other similar medical experiments? What is the connection. Is she trying to stir up sympathy for Henrietta and her family in a way that is not really justified? Do think she has slanted the story that way? If you really think this and can back it up, that would make for a stronger argument!

Astrid Ludwig

Content and Structure – The paper

1. (10) Has a clearly stated topic or question that it will address – 10

2. (5) Does address the stated topic or question – 5

3. Has a clearly articulated and logical structure:

a. (10) First paragraph lays out the topic or question – 10 b. (10) Later paragraphs have topic sentences and clearly identifiable main ideas –

9 c. (10) Transitions between paragraphs are logical – 9

4. (10) Accurately presents facts and viewpoints of sources – 8

5. (10) Provides sufficient textual evidence to support the argument – 10

6. (10) Is concise, clear, and not too repetitive – 9

7. (10) Has a clearly identifiable conclusion – 10

8. (15) Has a convincing argument that draws together the evidence presented – 8

Total: 88 (B+)

Style and English Usage – The paper

1. (20) Is clear, direct and appropriate in tone – 16

2. (10) Is free of misspellings – 10

3. (15) Is free of grammatical mistakes (e.g. subject/verb agreement, proper use of pronouns) – 12

4. (10) Lacks incomplete sentences – 10

5. (10) Lacks run-on or overly complicated sentences – 8

6. (10) Uses correct punctuation – 7

7. (15) Is free of jargon, cliches, and misused words – 13

8. (10) Cites quotations and/or references correctly – 10

Total: 85 (B)

Comments:

1. You say that “no one would care that much” if they knew who had originally donated the HeLa cell line. But what about Henrietta’s family? Didn’t they have some stake in this?

2. Your statements here are pretty certainly factually correct. But you should appreciate that there’s another side to this. As a matter of principle, couldn’t one say that a doctor should care about the human being behind a cell line that he or she uses? Isn’t that, in a way, one of the main points of Skloot’s book?? A person may or may not agree with that, of course. But if you want to argue that it’s not important or necessary for doctors and researchers to have that sort of compassion for the human beings whose tissues they are working with, then it seems to me that you have to make a case for that point of view.

3. Don’t you think Gey also had a certain desire to be remembered for donating the ”GeGe” cell line? Don’t forget, as a leader in the field before his death, his story certainly would be known to the later scientists who used those cells(!)

4. The idea of keeping the information a doctor gets from a patient confidential actually goes back a long way – it too is part of the original form of the Hippocratic Oath. The idea there was that no one would want to tell a doctor about their symptoms and other problems if they thought the doctor was going to go and reveal that information to others. So affirming confidentiality was “good business” for a doctor(!)

5-6. This is perhaps the most interesting aspect of this. What interference from the family

or future trouble do think Gey had in mind?

7. The sentence starting “The need for cells” is awkward and unclear. Please rewrite this to make your point more transparent.

Catherine Caracciolo

Content and Structure – The paper

1. (10 Has a clearly stated topic or question that it will address – 10

2. (5) Does address the stated topic or question – 5

3. Has a clearly articulated and logical structure:

a. (10) First paragraph lays out the topic or question – 10 b. (10) Later paragraphs have topic sentences and clearly identifiable main ideas –

9 c. (10) Transitions between paragraphs are logical – 9

4. (10) Accurately presents facts and viewpoints of sources – 8

5. (10) Provides sufficient textual evidence to support the argument – 10

6. (10) Is concise, clear, and not too repetitive – 9

7. (10) Has a clearly identifiable conclusion – 10

8. (15) Has a convincing argument that draws together the evidence presented – 8

Total: 88 (B+)

Style and English Usage – The paper

1. (20) Is clear, direct and appropriate in tone – 16

2. (10) Is free of misspellings – 9

3. (15) Is free of grammatical mistakes (e.g. subject/verb agreement, proper use of pronouns) – 12

4. (10) Lacks incomplete sentences – 10

5. (10) Lacks run-on or overly complicated sentences – 8

6. (10) Uses correct punctuation – 9

7. (15) Is free of jargon, cliches, and misused words – 11

8. (10) Cites quotations and/or references correctly – 10

Total: 85 (B)

Comments:

1. Can a case be ethical or not? Do you mean something like “How ethical was the treatment of Henrietta Lacks?”

2. It seems to me that the main point underlying everything you are saying is that you really think Skloot’s treatment of the Henrietta Lacks story is “slanted.” But you just throw that in as an assertion here without any real justification for it. I think that if you choose to rewrite the paper (and I would encourage you to do that), then that should be the main point you should argue, if you want to pursue the train of thought you started here. Your thesis might be that there is something wrong about the way Skloot combines Henrietta’s story with the stories of the Tuskegee syphilis study and the rest because the treatment she got at Hopkins is not comparable to the treatment that the victims of those other studies got. I’m not sure whether you need to mention the ”Unit 731” atrocities. Although they are similar in a way to things that are mentioned in the book, those particular “experiments” are not really relevant. Be careful, though. This is going to require some thought. Be sure you try to think about why Skloot chose to include these things before you say that what she did was slanted.

3. Is it correct to say that “no ethical codes were in place?”

4. Consider the source of this – does Jones have a vested interest?

5. Apart from the question of whether Henrietta’s treatment was standard for the time, isn’t it also reasonable to ask whether it followed what should have been the standard? Can’t we say that something was wrong in the past, even though it was common practice at the time? I think humans have the ability (at times even the responsibility), to come to the conclusion that some possible ethical standards are simply incorrect or indefensible.

6. You word choices (“riddled with ethical guidelines,” “engraved codes”) sound as though you think that higher ethical standards are not necessarily a good thing(!) If you don’t mean that, be more careful about how you say things. If you do mean that, this is something else that requires more justification!

7. Again, even if that is true, shouldn’t we also ask whether it should be true??

8. This is precisely the most horrific part of the Tuskegee story as far as I can see. If you have a treatment for the disease like syphilis, how is it conceivably ethically justifiable to allow the study to go on and let the subjects die gradually? Is any knowledge you get from the study worth enough to make that acceptable?

9-10. As I indicated above, I don’t see that these other examples add much to what you want to say. And please reconsider if you are trying to argue some version of “everyone was doing it, so it was OK by the standards of the time.”

11. You could also replace the “this city” by “[Baltimore]”

12. Be careful. This is a quote attributed to Henrietta by Laure Aurelian, who heard it second hand from George Gey. If you think about it, there’s not much evidence that Henrietta ever said anything like this at all. So making that your final word is a bit risky!

Daniel Finn

Content and Structure – The paper

1. (10 Has a clearly stated topic or question that it will address – 10

2. (5) Does address the stated topic or question – 5

3. Has a clearly articulated and logical structure:

a. (10) First paragraph lays out the topic or question – 8 b. (10) Later paragraphs have topic sentences and clearly identifiable main ideas –

9 c. (10) Transitions between paragraphs are logical – 9

4. (10) Accurately presents facts and viewpoints of sources – 8

5. (10) Provides sufficient textual evidence to support the argument – 10

6. (10) Is concise, clear, and not too repetitive – 8

7. (10) Has a clearly identifiable conclusion – 10

8. (15) Has a convincing argument that draws together the evidence presented – 10

Total: 87 (B+)

Style and English Usage – The paper

1. (20) Is clear, direct and appropriate in tone– 16

2. (10) Is free of misspellings – 9

3. (15) Is free of grammatical mistakes (e.g. subject/verb agreement, proper use of pronouns) – 12

4. (10) Lacks incomplete sentences – 10

5. (10) Lacks run-on or overly complicated sentences – 8

6. (10) Uses correct punctuation – 9

7. (15) Is free of jargon, cliches, and misused words – 11

8. (10) Cites quotations and/or references correctly – 10

Total: 85 (B)

Comments:

1. Don’t forget that Henrietta was also a rather uncooperative patient who had a longestablished pattern of avoiding follow-up treatment. So the fact that the cancer was not found earlier could also be explained by the fact she was not getting regular checkups and did not follow the doctors’ advice even when she did get medical attention.

2. It seems like an overstatement to say that these were the only forms of protection of rights of patients that were in place at the time. “Most important” might be closer to the mark.

3. Again, it seems like a big overstatement to say that there were few doctors who took ethics into account. I believe that Skloot’s points are that the ethical standards of the time for doctors were different from those accepted today (and questionable because of that), and also that our standards today are still possibly not ideal from the point of view of the rights of patients.

John Macomber

Content and Structure – The paper

1. (10 Has a clearly stated topic or question that it will address – 9

2. (5) Does address the stated topic or question – 5

3. Has a clearly articulated and logical structure:

a. (10) First paragraph lays out the topic or question – 9 b. (10) Later paragraphs have topic sentences and clearly identifiable main ideas –

9 c. (10) Transitions between paragraphs are logical – 9

4. (10) Accurately presents facts and viewpoints of sources – 7

5. (10) Provides sufficient textual evidence to support the argument – 5

6. (10) Is concise, clear, and not too repetitive – 9

7. (10) Has a clearly identifiable conclusion – 10

8. (15) Has a convincing argument that draws together the evidence presented – 12

Total: 84 (B)

Style and English Usage – The paper

1. (20) Is clear, direct and appropriate in tone – 16

2. (10) Is free of misspellings – 10

3. (15) Is free of grammatical mistakes (e.g. subject/verb agreement, proper use of pronouns) – 10

4. (10) Lacks incomplete sentences – 10

5. (10) Lacks run-on or overly complicated sentences – 10

6. (10) Uses correct punctuation – 10

7. (15) Is free of jargon, cliches, and misused words – 10

8. (10) Cites quotations and/or references correctly – 5

Total: 81 (B-)

Comments:

0. A general comment – there are many times in your paper where a quotation from the book would establish justification for a point you are making. But you have not done that. You have also not indicated where in the book the facts behind some of your claims are found. Please do this when you rewrite, and use the form given in the paper assignment.

1. I don’t think you want “the medical ethics raised” here. I think you mean something like “the treatment that Henrietta Lacks received.”

2. Don’t forget that it was not at all clear at the start that Henrietta Lacks had anything of “great value.” That only became clear after properties of the HeLa cell line became apparent.

3. “Radiation testing” is not correct here. Henrietta’s exposure to radium was a treatment designed to kill the cancer cells in her body.

4. As I said when we talked last week, I think the idea of making a comparison between the actions of the doctors in some of the situations described in the book and the actions of some of the people taking unwarranted risks on Wall Street in the recent crash might be interesting. But you need to develop this more. Exactly what about the situations is similar? Are there also differences?

5. Charles Southam injected cancer cells. There were rumors that some of the patients in the Tuskegee study were injected with syphilis, but I’m not sure that has ever been proved. Most of the subjects had been infected through sexual contacts and then entered the study where the disease was allowed to progress without any treatment.

Kaitlin O’Hagan

Content and Structure – The paper

1. (10) Has a clearly stated topic or question that it will address – 10

2. (5) Does address the stated topic or question – 5

3. Has a clearly articulated and logical structure:

a. (10) First paragraph lays out the topic or question – 9 b. (10) Later paragraphs have topic sentences and clearly identifiable main ideas –

9 c. (10) Transitions between paragraphs are logical – 9

4. (10) Accurately presents facts and viewpoints of sources – 10

5. (10) Provides sufficient textual evidence to support the argument – 10

6. (10) Is concise, clear, and not too repetitive – 9

7. (10) Has a clearly identifiable conclusion – 10

8. (15) Has a convincing argument that draws together the evidence presented – 7

Total: 88 (B+)

Style and English Usage – The paper

1. (20) Is clear, direct and appropriate in tone – 20

2. (10) Is free of misspellings – 10

3. (15) Is free of grammatical mistakes (e.g. subject/verb agreement, proper use of pronouns) – 15

4. (10) Lacks incomplete sentences – 10

5. (10) Lacks run-on or overly complicated sentences – 9

6. (10) Uses correct punctuation – 10

7. (15) Is free of jargon, cliches, and misused words – 15

8. (10) Cites quotations and/or references correctly – 10

Total: 99 (A)

Comments:

0. Your paper is very well-written. The major criticism I have is that you have really only considered one possible reason why the name “Henrietta Lacks” was kept secret so long – that George Gey and others might have wanted to monopolize profits from the HeLa cell line. As you show, that reason does not stand up to scrutiny. But it seems to me that there is more to the story than that. For instance, the fictitious names “Helen Lane,” etc. had to come from somewhere and there was some real deception involved in spreading them around. Why do you suppose that was done?

1. “Inconvenient” and “difficult” sound almost like excuses here. Couldn’t one argue that the doctors had a responsibility to try to make it clear to the family what they were going to do with the cells, even though it might have been difficult?

2. Why does it follow that they “had to conceal” Henrietta Lacks’s identity. I don’t see this.

3. “Juvenile selfishness” sounds pretty harsh! Don’t forget that in anything as difficult and time-consuming as scientific research, it is only natural for the people who make big advances and contributions to want to get proper credit for themselves!

Kevin Sutman

Content and Structure – The paper

1. (10) Has a clearly stated topic or question that it will address – 9

2. (5) Does address the stated topic or question – 5

3. Has a clearly articulated and logical structure:

a. (10) First paragraph lays out the topic or question – 8 b. (10) Later paragraphs have topic sentences and clearly identifiable main ideas –

9 c. (10) Transitions between paragraphs are logical – 9

4. (10) Accurately presents facts and viewpoints of sources – 10

5. (10) Provides sufficient textual evidence to support the argument – 7

6. (10) Is concise, clear, and not too repetitive – 9

7. (10) Has a clearly identifiable conclusion – 10

8. (15) Has a convincing argument that draws together the evidence presented – 7

Total: 83 (B/B-)

Style and English Usage – The paper

1. (20) Is clear, direct and appropriate in tone – 17

2. (10) Is free of misspellings – 10

3. (15) Is free of grammatical mistakes (e.g. subject/verb agreement, proper use of pronouns) – 10

4. (10) Lacks incomplete sentences – 10

5. (10) Lacks run-on or overly complicated sentences – 9

6. (10) Uses correct punctuation – 9

7. (15) Is free of jargon, cliches, and misused words – 13

8. (10) Cites quotations and/or references correctly – 8

Total: 86 (B)

Comments:

1. This is an overstatement and a misunderstanding of a key point in this story. There are certainly widely-accepted ethical principles for the practice of medicine that go back to the Hippocratic Oath you mention. There are also a number of federal and state laws in the US that regulate the practice of medicine. It is true that there are no legal requirements that doctors must get informed consent for experimental work on tissue taken from a patient’s body, which is the main point for the story of Henrietta Lacks. Note that that is a much more restricted issue.

2. Be careful about this – “Who told you you could sell my spleen?” is the title of a chapter in the book. It’s not clear from what’s in the book that John Moore ever said exactly that, and the book doesn’t claim that he did.

3. Again, be careful. The quotation you are using is actually Skloot reporting something like a headline from a newspaper story about the legal appeals in the Moore case. You shouldn’t use it this way.

4. Your last sentence does not really fit with what you were saying on the previous page about whether it was ethical to use Henrietta’s cells without consent. This could be clearer.

Kyle Kadlick

Content and Structure – The paper

1. (10) Has a clearly stated topic or question that it will address – 9

2. (5) Does address the stated topic or question – 5

3. Has a clearly articulated and logical structure:

a. (10) First paragraph lays out the topic or question – 8 b. (10) Later paragraphs have topic sentences and clearly identifiable main ideas –

9 c. (10) Transitions between paragraphs are logical – 7

4. (10) Accurately presents facts and viewpoints of sources – 10

5. (10) Provides sufficient textual evidence to support the argument – 9

6. (10) Is concise, clear, and not too repetitive – 8

7. (10) Has a clearly identifiable conclusion – 10

8. (15) Has a convincing argument that draws together the evidence presented – 11

Total: 86 (B)

Style and English Usage – The paper

1. (20) Is clear, direct and appropriate in tone – 13

2. (10) Is free of misspellings – 10

3. (15) Is free of grammatical mistakes (e.g. subject/verb agreement, proper use of pronouns) – 12

4. (10) Lacks incomplete sentences – 10

5. (10) Lacks run-on or overly complicated sentences – 9

6. (10) Uses correct punctuation – 8

7. (15) Is free of jargon, cliches, and misused words – 12

8. (10) Cites quotations and/or references correctly – 9

Total: 83 (B/B-)

Comments:

1. This is not a huge point, but in the paper assignment, I said you should say something like (Skloot, 38) for a reference for a direct quotation. The reason of course, is that if you had more than one source, you would need to say which one the quotation was drawn from.

2. This point does not really fit with the rest of the paragraph. You should make this into a paragraph on its own if you want to keep it. But on the other hand, what is the point you want to make with this? That Gey knew his job well enough to select assistants with the necessary skills? That seems like a relatively minor positive attribute to be singling it out like this.

3. In order to maintain Henrietta’s privacy, why was it necessary to be deceptive about the name of the subject? Did he have to lie to throw people off? Wouldn’t it have been enough just to say “HeLa came from a subject whose privacy I want to maintain, so I’m not going to tell you?” If you think about it, too, Henrietta was already dead by then, so why was her privacy such an issue any more?

4. Be careful. This is a quote attributed to Henrietta by Laure Aurelian, who heard it second hand from George Gey. If you think about it, there’s not much evidence that Henrietta ever said anything like this at all.

Lucia Bostrom

Content and Structure – The paper

1. (10) Has a clearly stated topic or question that it will address – 8

2. (5) Does address the stated topic or question – 5

3. Has a clearly articulated and logical structure:

a. (10) First paragraph lays out the topic or question – 8 b. (10) Later paragraphs have topic sentences and clearly identifiable main ideas –

9 c. (10) Transitions between paragraphs are logical – 9

4. (10) Accurately presents facts and viewpoints of sources – 8

5. (10) Provides sufficient textual evidence to support the argument – 9

6. (10) Is concise, clear, and not too repetitive – 8

7. (10) Has a clearly identifiable conclusion – 10

8. (15) Has a convincing argument that draws together the evidence presented – 12

Total: 86 (B)

Style and English Usage – The paper

1. (20) Is clear, direct and appropriate in tone – 15

2. (10) Is free of misspellings – 10

3. (15) Is free of grammatical mistakes (e.g. subject/verb agreement, proper use of pronouns) – 15

4. (10) Lacks incomplete sentences – 10

5. (10) Lacks run-on or overly complicated sentences – 9

6. (10) Uses correct punctuation – 10

7. (15) Is free of jargon, cliches, and misused words – 15

8. (10) Cites quotations and/or references correctly – 10

Total: 94 (A)

Comments:

1. Careful–the Nuremberg Trials were military tribunals held after WWII in which living former Nazis were accused and prosecuted as war criminals. The defendants in one series of trials were doctors who had been involved in “experimentation” on human subjects in the concentration camps.

2. True, but is that the whole story? Just because something is not required, couldn’t you still argue that it should be required?

3. Apart from the question of whether Henrietta’s treatment was standard for the time, isn’t it also reasonable to ask whether it followed what should have been the standard?

Can’t we say that something in the past was wrong, even though it was common practice at the time?

4. See point 1 above again. The Nuremberg Trials were not what you seem to think.

5. Do you really think that the definition of “ethics” is so ambiguous? Or is it that there are sometimes disagreements about what constitutes ethical or proper behavior?

Thomas Santa Maria

Content and Structure – The paper

1. (10) Has a clearly stated topic or question that it will address – 10

2. (5) Does address the stated topic or question – 5

3. Has a clearly articulated and logical structure:

a. (10) First paragraph lays out the topic or question – 9 b. (10) Later paragraphs have topic sentences and clearly identifiable main ideas –

10 c. (10) Transitions between paragraphs are logical – 10

4. (10) Accurately presents facts and viewpoints of sources – 10

5. (10) Provides sufficient textual evidence to support the argument – 10

6. (10) Is concise, clear, and not too repetitive – 10

7. (10) Has a clearly identifiable conclusion – 10

8. (15) Has a convincing argument that draws together the evidence presented – 12

Total: 96 (A)

Style and English Usage – The paper

1. (20) Is clear, direct and appropriate in tone – 20

2. (10) Is free of misspellings – 10

3. (15) Is free of grammatical mistakes (e.g. subject/verb agreement, proper use of pronouns) – 15

4. (10) Lacks incomplete sentences – 10

5. (10) Lacks run-on or overly complicated sentences – 9

6. (10) Uses correct punctuation – 9

7. (15) Is free of jargon, cliches, and misused words – 15

8. (10) Cites quotations and/or references correctly – 10

Total: 99 (A)

Comments:

1. When I read this again, it seemed to me now that (even though you say “in a graver sense”) the situation described in the opening of the first paragraph is so minor in comparison to the situations described in the book that saying “scenarios like this” just doesn’t fit. The difference in degree is too large for the comparison even to make sense. I think you really need a different example if you want to open with one.

2. The conclusion is not as strong as it could be. It sounds as though you are being careful, hedging your bets.

Muhammad Aqib Javaid

Content and Structure – The paper

1. (10) Has a clearly stated topic or question that it will address – 10

2. (5) Does address the stated topic or question – 5

3. Has a clearly articulated and logical structure:

a. (10) First paragraph lays out the topic or question – 8 b. (10) Later paragraphs have topic sentences and clearly identifiable main ideas –

9 c. (10) Transitions between paragraphs are logical – 8

4. (10) Accurately presents facts and viewpoints of sources – 10

5. (10) Provides sufficient textual evidence to support the argument – 10

6. (10) Is concise, clear, and not too repetitive – 10

7. (10) Has a clearly identifiable conclusion – 10

8. (15) Has a convincing argument that draws together the evidence presented – 13

Total: 93 (A-)

Style and English Usage – The paper

1. (20) Is clear, direct and appropriate in tone – 18

2. (10) Is free of misspellings – 10

3. (15) Is free of grammatical mistakes (e.g. subject/verb agreement, proper use of pronouns) – 11

4. (10) Lacks incomplete sentences – 10

5. (10) Lacks run-on or overly complicated sentences – 10

6. (10) Uses correct punctuation – 8

7. (15) Is free of jargon, cliches, and misused words – 15

8. (10) Cites quotations and/or references correctly – 10

Total: 92 (A-)

Comments:

0. There are some relatively minor technical problems here, but overall your writing is good. It is clear and forceful and there is no doubt at all what you think about the issues you raise. Beginning with the Southam quotation was an interesting choice, but I think it really worked to make your point about his self-serving and immoral attitudes.

1. On page 172, Skloot says that according to the experts she consulted, the fact that the tumor type was misdiagnosed did not change the treatment that Henrietta would have received at the time. Doctors did not have that many different treatment options at the time.

2. I think you could even ask whether it would be possible to have a proper justification for doing something like Southam’s cancer cell injection “experiment.”

Sarah Cahn

Content and Structure – The paper

1. (10) Has a clearly stated topic or question that it will address – 7

2. (5) Does address the stated topic or question – 5

3. Has a clearly articulated and logical structure:

a. (10) First paragraph lays out the topic or question – 10 b. (10) Later paragraphs have topic sentences and clearly identifiable main ideas –

10 c. (10) Transitions between paragraphs are logical – 10

4. (10) Accurately presents facts and viewpoints of sources – 7

5. (10) Provides sufficient textual evidence to support the argument – 9

6. (10) Is concise, clear, and not too repetitive – 8

7. (10) Has a clearly identifiable conclusion – 10

8. (15) Has a convincing argument that draws together the evidence presented – 5

Total: 81 (B-)

Style and English Usage – The paper

1. (20) Is clear, direct and appropriate in tone – 20

2. (10) Is free of misspellings – 10

3. (15) Is free of grammatical mistakes (e.g. subject/verb agreement, proper use of pronouns) – 14

4. (10) Lacks incomplete sentences – 10

5. (10) Lacks run-on or overly complicated sentences – 10

6. (10) Uses correct punctuation – 10

7. (15) Is free of jargon, cliches, and misused words – 10

8. (10) Cites quotations and/or references correctly – 10

Total: 99 (A)

Comments:

0. Your writing is quite strong technically. However I have some major issues with some of the points you wanted to make here.

1. A minor point first – Skloot’s book is definitely not a novel. It is not a work of fiction!

2. The first major point is one that I was trying to get you to see when you came to office hours to discuss the paper. First, I do not know what you mean by saying that “responsibility” is an ethical standard. An ethical standard is a general statement about what constitutes correct or moral behavior (perhaps in a particular situation). So you might say “a doctor should be responsible” as an ethical standard. But even that is pretty vague. Similarly, later when you discuss the Lacks family’s attitude, you might say their primary ethical standard was that “everyone should be treated fairly.” In order to say that what the doctors running the Tuskegee syphilis study did was ethical (according to some hypothetical standard), I think you would almost have to say they were acting according to a standard something like: “A doctor’s responsibility is to advance scientific knowledge and doing anything that advances scientific knowledge in whatever way is OK.” I don’t doubt that some people (possibly the Tuskegee doctors) set that sort of standard for themselves. But I hoped you would see how terribly wrong that would be – that makes it possible to inflict almost limitless cruelty on other people in the process! In other words, I think humans have the ability (at times even the responsibility), to come to the conclusion that some possible ethical standards are simply incorrect or indefensible.

3. In order to maintain Henrietta’s privacy, was it necessary to be deceptive about the name of the subject? Did Gey have to lie to throw people off? Wouldn’t it have been enough just to say “HeLa came from a subject whose privacy I want to maintain, so I’m not going to tell you?” If you think about it, too, Henrietta was already dead by then. So why was her privacy such an issue any more?

4. The idea of keeping the information a doctor gets from a patient confidential actually goes back a long way – it too is part of the original form of the Hippocratic Oath. The idea there was that no one would want to tell a doctor about their symptoms and other problems if they thought the doctor was going to go and reveal that information to others. So affirming confidentiality was “good business” for a doctor(!)

5. How could the word “benevolent” possibly be applied to anything connected with the Tuskegee study?? If you can’t exactly define the word, please look it up and then I hope you will understand why I was tearing my hair out when I read this!

6. The meaning of this sentence is not very clear since you have taken the quotation out of a larger context that dealt with proposed legislation. You need to explain that part of it if you want to use the “bills were repeatedly turned down ... ”

Peter Zona

Content and Structure – The paper

1. (10) Has a clearly stated topic or question that it will address – 8

2. (5) Does address the stated topic or question – 5

3. Has a clearly articulated and logical structure:

a. (10) First paragraph lays out the topic or question – 8 b. (10) Later paragraphs have topic sentences and clearly identifiable main ideas –

8 c. (10) Transitions between paragraphs are logical – 8

4. (10) Accurately presents facts and viewpoints of sources – 8

5. (10) Provides sufficient textual evidence to support the argument – 7

6. (10) Is concise, clear, and not too repetitive – 10

7. (10) Has a clearly identifiable conclusion – 10

8. (15) Has a convincing argument that draws together the evidence presented – 8

Total: 80 (B-)

Style and English Usage – The paper

1. (20) Is clear, direct and appropriate in tone – 20

2. (10) Is free of misspellings – 10

3. (15) Is free of grammatical mistakes (e.g. subject/verb agreement, proper use of pronouns) – 14

4. (10) Lacks incomplete sentences – 10

5. (10) Lacks run-on or overly complicated sentences – 10

6. (10) Uses correct punctuation – 10

7. (15) Is free of jargon, cliches, and misused words – 10

8. (10) Cites quotations and/or references correctly – 10

Total: 99 (A)

Comments:

1. This is something that reasonable people might disagree about, especially regarding the issue of informed consent.

2. On page 172, Skloot says that according to the experts she consulted, the fact that the tumor type was misdiagnosed did not change the treatment that Henrietta would have received at the time. Doctors did not have that many different treatment options at the time and it is doubtful that any of them could have cured Henrietta’s illness

3. I think this might be valid if you could demonstrate that the amounts charged for the cells were way out of line with the costs of producing them. If not, then I do not see why charging for the cells is unethical.

4. In order to maintain Henrietta’s privacy, was it necessary to be deceptive about the name of the subject? Did Gey have to lie to throw people off? Wouldn’t it have been enough just to say “HeLa came from a subject whose privacy I want to maintain, so I’m not going to tell you?” If you think about it, too, Henrietta was already dead by then. So why was her privacy such an issue any more?

5. This sentence is not correct – it certainly was not Henrietta or her family who were responsible for the medical advances. They were not doing the research. You could say Henrietta made those advances possible in a way, though.

Matt Kasuba

Content and Structure – The paper

1. (10) Has a clearly stated topic or question that it will address – 8

2. (5) Does address the stated topic or question – 5

3. Has a clearly articulated and logical structure:

a. (10) First paragraph lays out the topic or question – 8 b. (10) Later paragraphs have topic sentences and clearly identifiable main ideas –

9 c. (10) Transitions between paragraphs are logical – 9

4. (10) Accurately presents facts and viewpoints of sources – 10

5. (10) Provides sufficient textual evidence to support the argument – 8

6. (10) Is concise, clear, and not too repetitive – 10

7. (10) Has a clearly identifiable conclusion – 10

8. (15) Has a convincing argument that draws together the evidence presented – 8

Total: 85 (B)

Style and English Usage – The paper

1. (20) Is clear, direct and appropriate in tone – 18

2. (10) Is free of misspellings – 10

3. (15) Is free of grammatical mistakes (e.g. subject/verb agreement, proper use of pronouns) – 13

4. (10) Lacks incomplete sentences – 10

5. (10) Lacks run-on or overly complicated sentences – 10

6. (10) Uses correct punctuation – 10

7. (15) Is free of jargon, cliches, and misused words – 11

8. (10) Cites quotations and/or references correctly – 10

Total: 92 (A-)

Comments:

1. This is an overstatement. There were certainly generally-accepted ethical principles about aspects of all of this that doctors followed at the time. If you mean “code” in the sense of a set of legal restrictions, that is something different And as Skloot points out, there are still no legal bar to doctors using tissue taken from a patient’s body for experimentation.

2. This quotation is a very troubling one, I think, because it shows the way racial prejudice contributed to the treatment of the subjects in the Tuskegee study. It is almost as if the doctors carrying out the study could justify what they were doing by falling back on the idea that African Americans were inferior, maybe even not quite fully human.”

Max Pettinelli

Content and Structure – The paper

1. (10) Has a clearly stated topic or question that it will address – 10

2. (5) Does address the stated topic or question – 5

3. Has a clearly articulated and logical structure:

a. (10) First paragraph lays out the topic or question – 10 b. (10) Later paragraphs have topic sentences and clearly identifiable main ideas –

9 c. (10) Transitions between paragraphs are logical – 9

4. (10) Accurately presents facts and viewpoints of sources – 10

5. (10) Provides sufficient textual evidence to support the argument – 8

6. (10) Is concise, clear, and not too repetitive – 10

7. (10) Has a clearly identifiable conclusion – 10

8. (15) Has a convincing argument that draws together the evidence presented – 12

Total: 93 (A-)

Style and English Usage – The paper

1. (20) Is clear, direct and appropriate in tone – 18

2. (10) Is free of misspellings – 10

3. (15) Is free of grammatical mistakes (e.g. subject/verb agreement, proper use of pronouns) – 11

4. (10) Lacks incomplete sentences – 10

5. (10) Lacks run-on or overly complicated sentences – 10

6. (10) Uses correct punctuation – 10

7. (15) Is free of jargon, cliches, and misused words – 11

8. (10) Cites quotations and/or references correctly – 10

Total: 94 (A)

Comments:

1. I think it is doubtful that Henrietta’s cancer was in part caused by the radium treatments. The interaction of her cells with the HPV virus seems to be the accepted explanation for the properties of her cells.

2. Careful – the ethical issue is whether doctors should have to get consent for using tissues collected from a patient, not the collection of the tissue itself.

3. Why is it so clear that patients should not sell bits of themselves? Isn’t that something that could be left as an individual choice? In other words, if someone like John Moore wanted to give his consent to have his spleen used for research in return for some monetary consideration, why that be improper?

Xuan Ha Nghiem

Content and Structure – The paper

1. (10) Has a clearly stated topic or question that it will address – 10

2. (5) Does address the stated topic or question – 5

3. Has a clearly articulated and logical structure:

a. (10) First paragraph lays out the topic or question – 10 b. (10) Later paragraphs have topic sentences and clearly identifiable main ideas –

9 c. (10) Transitions between paragraphs are logical – 9

4. (10) Accurately presents facts and viewpoints of sources – 10

5. (10) Provides sufficient textual evidence to support the argument – 10

6. (10) Is concise, clear, and not too repetitive – 10

7. (10) Has a clearly identifiable conclusion – 10

8. (15) Has a convincing argument that draws together the evidence presented – 15

Total: 98 (A)

Style and English Usage – The paper

1. (20) Is clear, direct and appropriate in tone – 18

2. (10) Is free of misspellings – 10

3. (15) Is free of grammatical mistakes (e.g. subject/verb agreement, proper use of pronouns) – 10

4. (10) Lacks incomplete sentences – 10

5. (10) Lacks run-on or overly complicated sentences – 8

6. (10) Uses correct punctuation – 9

7. (15) Is free of jargon, cliches, and misused words – 15

8. (10) Cites quotations and/or references correctly – 10

Total: 90 (A-)

Comments:

You have a lot of good ideas about issues raised in Skloot’s book. Your writing is good, but there are a few sentences that still need some polishing. I have been rather “picky” again in the comments on the paper. Look these over and let me know if you have any questions.

Kylee Sullivan

Content and Structure – The paper

1. (10) Has a clearly stated topic or question that it will address – 8

2. (5) Does address the stated topic or question – 5

3. Has a clearly articulated and logical structure:

a. (10) First paragraph lays out the topic or question – 9 b. (10) Later paragraphs have topic sentences and clearly identifiable main ideas –

9 c. (10) Transitions between paragraphs are logical – 9

4. (10) Accurately presents facts and viewpoints of sources – 9

5. (10) Provides sufficient textual evidence to support the argument – 10

6. (10) Is concise, clear, and not too repetitive – 10

7. (10) Has a clearly identifiable conclusion – 10

8. (15) Has a convincing argument that draws together the evidence presented – 8

Total: 87 (B+)

Style and English Usage – The paper

1. (20) Is clear, direct and appropriate in tone – 18

2. (10) Is free of misspellings – 10

3. (15) Is free of grammatical mistakes (e.g. subject/verb agreement, proper use of pronouns) – 15

4. (10) Lacks incomplete sentences – 10

5. (10) Lacks run-on or overly complicated sentences – 8

6. (10) Uses correct punctuation – 9

7. (15) Is free of jargon, cliches, and misused words – 15

8. (10) Cites quotations and/or references correctly – 9

Total: 96 (A)

Comments:

1. I think this is too general and overstated. There is plenty of unethical behavior to go around in almost every era of human history and singling out the 1950’s as especially “unethical” does not seem justified. Plus, what does it mean for a time to be ethical or unethical, anyway? I think you mean the prevailing ethical standards in medical practice seem questionable today. Also, at the end of the first paragraph you say that the standards of the 1950’s were “sensible and justifiable” at that time. But it seems to me you need to make distinctions here–it’s not possible to make blanket statements like this that cover everything about a subject as complicated as this for a whole era. You do, in fact, draw some distinctions later in the paper. So I think you recognized this. The introductory paragraph should acknowledge that. Moreover, I think it would be good to try to say why or how people could justify medical experimentation without full consent. What arguments would they make to say that was justified? Are there problems with those arguments?

2. Are you saying that the fact that racial discrimination was law at the time (a true statement, of course) somehow made it OK? If not, then be more careful about how you say this. If so, can’t we say that present or past laws are unjust by standards to which we aim to hold ourselves?

3. “Colored” is a word that has too many negative associations (because of the racial prejudice of the time when it was commonly used). It’s better to stick to “African American” unless you are reporting direct speech from the time.

4. Charles Southam injected cancer cells. There were rumors that some of the patients in the Tuskegee study were injected with syphilis, but I’m not sure that has ever been proved. Most of the subjects had been infected through sexual contacts and then entered the study where the disease was allowed to progress without any treatment.

5. This is a place where, even though you are not using a direct quotation, it would still be good to give a reference to the discussion in the book about the first woman that Southam injected with the HeLa cells.

6. I don’t think I really follow your concluding paragraph. You say that codes of ethics are “objective,” but then most of what follows talks about how standards can change over time. “Objective” seems to indicate that there are clearcut principles of what is right and what is wrong that don’t change. So I hope you understand my confusion. I think this is the ultimate question you need to think about here: Don’t humans have the ability (at times even the responsibility), to come to the conclusion that some possible ethical standards are simply incorrect or indefensible.